OpenStudy (anonymous):

Find g^-1(x) when g(x)=sqrt(x+3)

OpenStudy (anonymous):

\[y=\sqrt{x+3}\] first you interchane the variable x and y \[x=\sqrt{y+3}\] solve for y by squaring both sides \[x^2=y+3\] subtract 3 from both sides to get \[x^2-3=y\]

OpenStudy (anonymous):

I see! =) So my final answer should be y=x^2-3

OpenStudy (anonymous):

yep there is a systematic way of doing this 1)interchange x and y 2)solve for y

OpenStudy (anonymous):

Thank you so much!

OpenStudy (anonymous):

there is a way to check this also

OpenStudy (anonymous):

since we know that the domain of sqrt of x+3 is as semi circle with radius 3 , we can see that they're inverses that are reflected upon the x axis

OpenStudy (anonymous):